### 06 Apr 2026 v1

@@ 5 Complexity analysis
First, we estimate the number of inner iterations, denoted by $N_{\text{in}}$. Fix an outer iteration index $k$ and suppose that $\xi_{\rho}(\hat{x}^{(k)},s^{(k)};\mu^{(k)})\leq\kappa$. For a fixed point $w^{(k)}=(\hat{x}^{(k)},s^{(k)})$ and $\rho\geq 1$, define the merit function $\psi:\mathbb{R}\times\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ as
$\psi(\mu,h):=\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}}{\mu S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}.$ (128)
A direct computation gives
$\psi^{\prime}(\mu,h)$ $=\frac{2\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle\cdot\langle\nabla_{w}\eta_{\rho}^{\prime}(w^{(k)};\mu),h\rangle}{\mu S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}-\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}}{\mu^{2}S_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}$ (129)
$\qquad\qquad-\frac{\langle\nabla_{w}\eta_{\rho}(w^{(k)};\mu),h\rangle^{2}S^{\prime}_{\eta_{\rho}}(w^{(k)};\mu)[h,h]}{\mu(S_{\eta_{\rho}}(w^{(k)};\mu)[h,h])^{2}}.$
Let $\mu\in[\mu^{(k+1)},\mu^{(k)}]$. It follows from Theorems 5.25.3 that
$|\psi^{\prime}(\mu,h)|$ $\leq\sqrt{\frac{8\nu}{\mu^{2}}}\sqrt{\psi(\mu,h)}+\frac{1+\rho(1+2\sqrt{\nu})}{\mu}\psi(\mu,h)$ (130)
$\leq\sqrt{\frac{8\nu}{(\mu^{(k+1)})^{2}}}\sqrt{\psi(\mu,h)}+\frac{1+\rho(1+2\sqrt{\nu})}{\mu^{(k+1)}}\psi(\mu,h).$
Define $c_{1}:=\frac{1+\rho(1+2\sqrt{\nu})}{2\mu^{(k+1)}}$ and $\Psi(\mu,h):=e^{c_{1}\mu}\sqrt{\psi(\mu,h)}$. We have
$\Psi^{\prime}(\mu,h)\leq e^{c_{1}\mu}\frac{\sqrt{2\nu}}{\mu^{(k+1)}}.$ (131)
Integrating both sides of (131) over $\mu^{(k+1)},\mu^{(k)}]$ yields
$\Psi(\mu^{(k+1)},h)-\Psi(\mu^{(k)},h)\leq\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}\mu^{(k)}}-e^{c_{1}\mu^{(k+1)}}\right).$ (132)
$\sqrt{\psi(\mu^{(k+1)},h)}$ (133)
$\leq$ $\,e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\sqrt{\psi(\mu^{(k)},h)}+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right)$
$\leq$ $\,e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\xi_{\rho}(\hat{x},s;\mu^{(k)})+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right),$
$\xi_{\rho}(\hat{x},s;\mu^{(k+1)})$ $\leq e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}\xi_{\rho}(\hat{x},s;\mu^{(k)})$ (134)
$\qquad\qquad\qquad+\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}-1\right).$
Recall $\mu^{(k+1)}=\sigma\mu^{(k)}$. Let
$\left\{\begin{array}[]{ll}&\alpha_{1}(\sigma,\rho):=e^{c_{1}(\mu^{(k)}-\mu^{(k+1)})}=e^{\frac{\rho(1+2\sqrt{\nu})+1}{2}(\frac{1}{\sigma}-1)},\\ &\alpha_{2}(\rho):=\frac{\sqrt{2\nu}}{\mu^{(k+1)}}\frac{1}{c_{1}}=\frac{2\sqrt{2\nu}}{\rho(1+2\sqrt{\nu})+1}.\end{array}\right$. (135)
It can be verified that
$\alpha_{1}(\sigma,\rho)\leq e^{\frac{\rho(1+2\sqrt{\nu})+1}{2}\cdot\frac{\ln(\gamma)}{2\rho\sqrt{\nu}}}\leq\gamma\text{ \, and\, }\alpha_{2}(\rho)\leq\sqrt{2}/\rho,$ (136)
whenever $\rho\geq 1$ and $1>\sigma\geq 1-\dfrac{\ln(\gamma)}{2\rho\sqrt{\nu}+\ln(\gamma)}$. At the end of $k$-th iteration, one has $\xi_{\rho}(w^{(k)};\mu^{(k)})\leq\kappa$. It follows immediately that
$\xi_{\rho}(w^{(k)};\mu^{(k+1)})\leq\alpha_{1}(\sigma,\rho)\kappa+\alpha_{2}(\rho)(\alpha_{1}(\sigma,\rho)-1)\leq 2\kappa.$ (137)
By Theorem 2.3(iii), one Newton step suffices to obtain
$\xi_{\rho}(\hat{x}^{(k+1)},s^{(k+1)};\mu^{(k+1)})\leq\kappa.$
Hence, $N_{\rm in}=1$.
Next, we estimate the number of outer iterations, denoted by $N_{\text{out}}$. Since $\mu^{(k)}=\sigma^{k}\mu^{(0)}$, termination of Algorithm 2 at $k$-th iteration implies
$\mu^{(k)}\leq\varepsilon\quad\Longrightarrow\quad k\geq\frac{\ln({\mu^{(0)}}/{\varepsilon})}{\ln(1/\sigma)}.$ (138)
Consequently,
$N_{\text{out}}\leq-\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)/\ln(\sigma)+1=\mathcal{O}\!\left(\sqrt{\nu}\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right).$ (139)
The total number of iterations is given by
$N_{\text{in}}\times N_{\text{out}}=\mathcal{O}\left(\sqrt{\nu}\ln\left(\dfrac{\mu^{(0)}}{\varepsilon}\right)\right),$
which completes the proof.
@@
Combining Theorems 5.1 and 5.5, we obtain that the first-phase of PFSNM requires at most $\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ iterations, whereas the second-phase of PFSNM admits an iteration complexity of $\mathcal{O}\left(\sqrt{\nu}\ln\left(\mu^{(0)}/{\varepsilon}\right)\right)$. Therefore, the overall iteration complexity is
$\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)+\mathcal{O}\left(\sqrt{\nu}\ln\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right).$
Since $\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}$ is independent of $\varepsilon$, the above bound can be written more compactly as $\mathcal{O}\left(\sqrt{\nu}\ln\left(1/{\varepsilon}\right)\right)$.
@@ 6 Computational results
We evaluate solver performance employing performance profiles Dolan and Moré (2002) and the shifted geometric mean [^4] https://plato.asu.edu/ftp/shgeom.html (SGM). Let $\mathcal{P}$ denote the benchmark set and $\mathcal{S}$ the set of solvers. For each problem $p\in\mathcal{P}$ and solver $s\in\mathcal{S}$, let $t_{p,s}$ be the runtime, and define the performance ratio
$r_{p,s}=\frac{t_{p,s}}{\min_{s^{\prime}\in\mathcal{S}}t_{p,s^{\prime}}}\in[1,\infty],$ (140)
@@
$\rho_{s}(\tau)=\frac{1}{|\mathcal{P}|}\Big|\big\{p\in\mathcal{P}:\ r_{p,s}\leq\tau\big\}\Big|,\qquad\tau\geq 1,$ (141)
@@
$\mathrm{SGM}_{s}=\exp\!\left(\sum_{p\in\mathcal{P}}\ln\left(\frac{\max\{1,t_{p,s}+\text{offset}\}}{|\mathcal{P}|}\right)\right)-\text{offset},$ (142)
so that smaller values indicate better aggregate efficiency while remaining insensitive to a small number of difficult instances.
@@ 6.1 Linear programs
We first test the solvers on linear programs from the widely used NETLIB collection. Each solver is run with the same time limit and accuracy requirements, and the results are summarized by Table 2 and Fig. 3.
Table 2 reports the solved-problem ratio together with the shifted geometric mean. The main observation is that PFSNM achieves the strongest overall combination of robustness and efficiency on this benchmark. In particular, it achieves the best aggregate runtime behavior according to the SGM, and it does so without compromising reliability, whereas competing solvers exhibit either a lower success rate or a noticeably larger SGM. This indicates that the advantage of PFSNM on NETLIB is consistent across instances, reflecting a favorable overall balance between convergence robustness and computational cost.
Table 2: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection.
@@
Fig. 3 provides a more intuitive perspective through performance profiles. The curve of PFSNM stays above the competing solvers over essentially the entire range of $\tau$. This behavior indicates that PFSNM is frequently the fastest solver among the five solvers on a large portion of the benchmark set. Furthermore, $\rho_{s}(\tau)$ for PFSNM approaches 11 as $\tau$ increases, which means that it successfully solves all LP instances under the imposed limits. In contrast, the profiles of the other solvers level off below 11, and their slower rise for small $\tau$ indicates weaker competitiveness on instances with comparable runtimes.
Figure 3: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection.
@@ 6.2 Quadratic programs
Table 3 suggests that PFSNM is competitive in terms of aggregate efficiency, although it is not the top-performing solver overall. In particular, its shifted geometric mean is the second-best among the tested solvers (about $2.66$), whereas Clarabel attains the best value (normalized to $1.00$). At the same time, the solved-problem ratios indicate that Clarabel succeeds on a larger portion of the QP instances (about $91.3\%$), while PFSNM solves a smaller but still sizable fraction (about $82.6\%$). Overall, the table indicates that the main difference between PFSNM and the best solver on this benchmark lies in robustness on a subset of instances.
Table 3: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection.
@@
Fig. 4 provides a consistent view. The performance profile of PFSNM rises quickly for small values of $\tau$, which indicates that when PFSNM succeeds, it often achieves runtimes close to the best solver on those instances. However, the limiting value of its profile remains below that of the most reliable solver, reflecting the gap in solved ratios reported in Table 3. Overall, the QP results show that PFSNM can be fast on the instances it solves, while improving robustness on the harder subset of Maros–Mészáros problems would further enhance its performance on this benchmark.
Figure 4: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection.
@@ 6.3 Second-order cone programs
$\min\limits_{y\in\mathbb{R}^{n}}\left\{\|Dy-b\|_{2}+\varrho\|y\|_{1}\right\},$ (143)
where $D\in\mathbb{R}^{d\times n}$ is a matrix from the SuiteSparse Matrix Collection, $b\in\mathbb{R}^{d}$ is a given vector, and $\varrho$ is a penalty parameter. This problem is equivalent to the following SOCP instance:
$\min\left\{t+\varrho\sum\limits_{i=1}^{n}(y^{+}_{i}+y^{-}_{i})\,\Big|\,Dy^{+}-Dy^{-}-d=b,\,(t,d)\in\mathbb{Q}^{d+1},\,y^{+},y^{-}\in\mathbb{R}^{n}_{+}\right\}$ (144)
We follow the recent work Goulart and Chen (2024) and choose $\varrho=\|D^{\top}b\|_{\infty}$. The vector $b$ is set to be the all-ones vector. The results are reported in Table 4 and Fig. 5.
Table 4 shows that all solvers solve all the instances, so the comparison is driven by efficiency rather than robustness. Under this setting, PFSNM attains the smallest SGM (normalized to $1.0000$). The closest competitor is Clarabel, with a SGM only slightly larger (about $1.08$), whereas the remaining solvers have clearly larger SGM values. Hence, the table indicates that PFSNM provides the best aggregate runtime performance on these Lasso-type SOCPs, with a particularly tight competition against Clarabel.
Table 4: SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices.
@@
Fig. 5 corroborates the table-based summary. Since all solvers succeed, the key difference lies in how quickly each profile rises near $\tau=1$. The PFSNM curve increases the fastest and stays close to the best observed curve over a wide range of $\tau$, meaning that it achieves the best or near-best runtime on a large fraction of instances. Combined with the SGM results, these experiments indicate that PFSNM offers strong and reliable performance on SOCP problems constructed from SuiteSparse matrices.
Figure 5: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices.
@@ 7 Conclusion
The PFSNM has been proposed for SCP based on a reduced SBAL function. Its associated parameterized smooth system has been shown to be equivalent to the first-order optimality conditions of a structured minimax problem. This characterization makes it possible to analyze the method within a self-concordant convex-concave framework adapted to the reduced formulation. It has been proved that the reduced SBAL function is $\mu$-self-concordant convex-concave and that the resulting method attains a worst-case iteration complexity of $O(\sqrt{\nu}\ln(1/\varepsilon))$. This iteration complexity matches the best-known short-step bound for IPMs on symmetric cones. Moreover, the reduced formulation also yields Newton systems with an explicit Schur complement, which lowers the cost of system formation relative to existing smoothing Newton methods. Numerical results indicate that the method is competitive on standard conic benchmarks.

### 21 May 2026 v2

@@ 5 Complexity analysis
We first show the iterates remain in the neighborhood $\mathcal{N}(\kappa,\mu^{(k)},\rho)$ for any $k\geq 0$. The claim holds for $k=0$ by the initialization. Suppose that $v^{(k)}\in\mathcal{N}(\kappa,\mu^{(k)},\rho),$ which implies that
$\xi_{\rho}(w^{(k)};\mu^{(k)})\leq\kappa.$
If $\mu^{(k)}\leq\varepsilon$, the algorithm terminates. Otherwise, the smoothing parameter is reduced according to $\mu^{(k+1)}=\sigma\mu^{(k)}.$ Applying Lemma 4 with $w=w^{(k)},\,\mu=\mu^{(k)},\,\mu^{+}=\mu^{(k+1)},$ we obtain
$\xi_{\rho}(w^{(k)};\mu^{(k+1)})\leq 2\kappa.$
Since $\kappa=0.1$, we have $2\kappa=0.2<2-\sqrt{3}$. Let $w^{(k+1)}$ denote the point obtained after one full Newton step. By Theorem 2.3(iii)
$\xi_{\rho}(w^{(k+1)};\mu^{(k+1)})\leq\frac{\xi_{\rho}(w^{(k)};\mu^{(k+1)})}{2}\leq\kappa.$
This, combined with (65), implies that $v^{(k+1)}\in\mathcal{N}(\kappa,\mu^{(k+1)},\rho)$. By induction, we conclude that $v^{(k)}$ remains in the neighborhood $\mathcal{N}(\kappa,\mu^{(k)},\rho)$ for any $k\geq 0$.
Now we bound the number of Newton steps. Since $\mu^{(k)}=\sigma^{k}\mu^{(0)}$, the stopping criterion $\mu^{(k)}\leq\varepsilon$ of Algorithm 2 is satisfied whenever
$k\geq\frac{\ln({\mu^{(0)}}/{\varepsilon})}{\ln(1/\sigma)}.$
Consequently, Algorithm 2 requires at most
$\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)/\ln(\sigma)+1=\mathcal{O}\!\left(\sqrt{\nu}\ln\!\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right)$
iterations to obtain an approximate KKT triple.
@@
Combining Theorems 5.1 and 5.5, we obtain that the first phase of PFSNM requires at most $\mathcal{O}\left(1+\frac{4t^{(0)}}{\kappa}\Theta_{1}(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}})\right)$ iterations, whereas the second phase of PFSNM admits an iteration complexity of $\mathcal{O}\left(\sqrt{\nu}\ln\left(\mu^{(0)}/{\varepsilon}\right)\right)$. Therefore, the overall iteration complexity is
$\mathcal{O}\left(1+\frac{4t^{(0)}}{\kappa}\Theta_{1}\left(\dfrac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\right)+\mathcal{O}\left(\sqrt{\nu}\ln\left(\frac{\mu^{(0)}}{\varepsilon}\right)\right).$
Since the first term is independent of $\varepsilon$, the above bound can be written more compactly as $\mathcal{O}\left(\sqrt{\nu}\ln\left(1/{\varepsilon}\right)\right)$.
@@ 6 Computational results
We evaluate solver performance using performance profiles Dolan and Moré (2002) and the shifted geometric mean (SGM) [^4] https://plato.asu.edu/ftp/shgeom.html. Let $\mathcal{P}$ denote the benchmark set and $\mathcal{S}$ denote the set of solvers. For each problem $p\in\mathcal{P}$ and solver $s\in\mathcal{S}$, let $t_{p,s}$ be the runtime, and define the performance ratio
$r_{p,s}=\frac{t_{p,s}}{\min_{s^{\prime}\in\mathcal{S}}t_{p,s^{\prime}}}\in[1,\infty],$
@@
$\rho_{s}(\tau)=\frac{1}{|\mathcal{P}|}\Big|\big\{p\in\mathcal{P}:\ r_{p,s}\leq\tau\big\}\Big|,\qquad\tau\geq 1,$
@@
$\mathrm{SGM}_{s}=\exp\!\left(\sum_{p\in\mathcal{P}}\frac{1}{{|\mathcal{P}|}}\ln\left({\max\{1,t_{p,s}+\text{offset}\}}\right)\right)-\text{offset},$
so that smaller values indicate better aggregate efficiency while being relatively insensitive to a small number of difficult instances.
@@ 6.1 Linear programs
We first test the solvers on linear programs from the widely used NETLIB collection. The results are summarized in Table 2 and Fig. 3.
Table 2 reports the solved ratios together with the shifted geometric means. The main observation is that PFSNM achieves a favorable balance between robustness and efficiency on this benchmark. In particular, it achieves the best aggregate runtime behavior as measured by the SGM without compromising reliability, whereas competing solvers exhibit either a lower success rate or a noticeably larger SGM. This indicates that the advantage of PFSNM on NETLIB is consistent across instances, reflecting a favorable overall balance between convergence robustness and computational cost.
Table 2: Solved ratios and SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection.
@@
Figure 3 provides a more intuitive perspective through performance profiles. The curve of PFSNM stays above the competing solvers over essentially the entire range of $\tau$. This behavior indicates that PFSNM is frequently the fastest among the five solvers on a large fraction of the benchmark set. Furthermore, $\rho_{s}(\tau)$ for PFSNM approaches $1$ as $\tau$ increases, which means that it successfully solves all LP instances under the imposed limits. In contrast, the profiles of the other solvers level off below $1$, and their slower rise for small $\tau$ indicates weaker performance on instances with comparable runtimes.
Figure 3: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the NETLIB collection
@@ 6.2 Quadratic programs
Table 3 suggests that PFSNM is competitive in terms of aggregate efficiency, although it is not the best-performing solver overall. In particular, its shifted geometric mean is the second-best among the tested solvers (about $2.66$), whereas Clarabel attains the best value (normalized to $1.00$). At the same time, the solved ratios indicate that Clarabel succeeds on a larger fraction of the QP instances (about $91.3\%$), while PFSNM solves a smaller but still substantial fraction (about $82.6\%$). Overall, the table indicates that the main difference between PFSNM and the best solver on this benchmark lies in robustness on a subset of instances.
Table 3: Solved ratios and SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection.
@@
Figure 4 provides a consistent view. The performance profile of PFSNM rises quickly for small values of $\tau$, which indicates that when PFSNM succeeds, its runtimes are often close to those of the best solver on those instances. However, the limiting value of its profile remains below that of the most reliable solver, reflecting the gap in solved ratios reported in Table 3. Overall, the QP results show that PFSNM can be fast on the instances it solves. Improving robustness on the more difficult subset of Maros–Mészáros problems would further enhance its performance on this benchmark.
Figure 4: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on the Maros–Mészáros collection
@@ 6.3 Second-order cone programs
$\min\limits_{y\in\mathbb{R}^{n}}\left\{\|Dy-b\|_{2}+\varrho\|y\|_{1}\right\},$
where $D\in\mathbb{R}^{d\times n}$ is a matrix from the SuiteSparse Matrix Collection, $b\in\mathbb{R}^{d}$ is a given vector, and $\varrho$ is a penalty parameter. This problem is equivalent to the following SOCP reformulation:
$\min\left\{t+\varrho\sum\limits_{i=1}^{n}(y^{+}_{i}+y^{-}_{i})\,\Big|\,Dy^{+}-Dy^{-}-r=b,\,(t,r)\in\mathbb{Q}^{d+1},\,y^{+},y^{-}\in\mathbb{R}^{n}_{+}\right\}.$
Following Goulart and Chen (2024), we choose $\varrho=\|D^{\top}b\|_{\infty}$. The vector $b$ is set to the all-ones vector. The results are reported in Table 4 and Fig. 5.
Table 4 shows that all solvers solve all the instances, so the comparison is driven by efficiency rather than robustness. In this setting, PFSNM attains the smallest SGM (normalized to $1.0000$). The closest competitor is Clarabel, with an SGM only slightly larger (about $1.0818$), whereas the remaining solvers have clearly larger SGM values. Hence, the table indicates that PFSNM provides the best aggregate runtime performance on these Lasso-type SOCPs, with particularly close performance relative to Clarabel.
Table 4: Solved ratios and SGMs of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices.
@@
Figure 5 is consistent with the table-based summary. Since all solvers succeed, the key difference lies in how quickly each profile rises near $\tau=1$. The PFSNM curve increases the fastest and stays close to the best observed curve over a wide range of $\tau$, meaning that it achieves the best or near-best runtime on a large fraction of instances. Combined with the SGM results, these experiments indicate that PFSNM offers strong and reliable performance on SOCP problems constructed from SuiteSparse matrices.
Figure 5: Performance profiles of PFSNM, SDPT3, SeDuMi, ECOS, and Clarabel on SOCP problems constructed from SuiteSparse matrices
@@ 7 Conclusion
We have proposed a path-following smoothing Newton method for symmetric cone programming based on a reduced BAL function. The associated parameterized smooth system has been shown to be equivalent to the first-order optimality conditions of a structured minimax problem. This characterization makes it possible to analyze the proposed method within a self-concordant convex-concave framework. We have proved that the reduced BAL function is $\mu$-self-concordant convex-concave and that the resulting method attains a worst-case iteration complexity of $O(\sqrt{\nu}\ln(1/\varepsilon))$. This iteration complexity matches the best-known short-step bound for IPMs on symmetric cones. Moreover, the reduced BAL function also yields Newton systems with an explicit Schur complement, which can be exploited to reduce system-formation costs for important classes of symmetric cones. Numerical results indicate that the method is competitive on standard conic benchmarks.